# Distance Recomputator and Topology Reconstructor for Graph Neural Networks

### v1

@@ 4.1 The In-Depth Learning of Graph Neural Networks
for n in $V$ do
       for $k=K...1$ do
             $SN=RS(N,Layer^{(k-1)})$ for $v_{i}$ in layer(n,k-1) do
                   $NGH[k][n]\leftarrow NGH[k][n]\cup GraphSAGE(v_{i},SN[i])$
@@
$\mathcal{B^{K}}\leftarrow\mathcal{B}$; // The $\mathcal{B}$ denotes the current batch of nodes to be processed.for $k=K...1$ do
       $B^{k-1}\leftarrow B^{k}$; for $u_{i}\in\mathcal{B}$ do
             $SN_{s}\leftarrow RS(u,N)$$NGH[k][n]\leftarrow NGH[k][n]\cup GraphSAGE(u_{i},SN_{s}[i])$
            $z_{u_{i}}\leftarrow GKHDA(u_{i},k,NGH[k][n])$
@@
Algorithm 1 Preprocess Static Sampling and Dynamic Resampling Strategies
Input : Graph $\mathcal{G(V,E)}$;
input features: $x_{v}$
depth K;
non-linearity: $\sigma$;
differentiable aggregator functions $AGGREGATE_{k}$, $\forall k\in\{1,...,K\}$;
neighborhood sampling functions, $N_{k}:v\longrightarrow 2^{v}$, $\forall k\in\{1,...,K\}$;
NGH = $\leftarrow$ StaticSampling($\mathcal{G}$,K,Z,A);
weight matrices $W^{k}$; //learnable matrix
Output : Vector Representation $z_{v}$ for all $v\in\mathcal{B}$
$h_{u}^{0}\leftarrow x_{v}$, $\forall v\in B$ for $k=1...K$ do
       for $u\in B^{k}$ do
             $h^{k}_{u}\leftarrow\sigma(W^{k}\cdot(h_{u}^{k-1},h_{N(u)}^{k}))$; $h^{k}_{u}\leftarrow h^{k}_{u}/\lVert h^{k}_{u}\rVert_{2}$;
@@
for $k=K...1$ do
       for $u\in NGH[u,k]$ do
             $h^{k}_{N(u)}\leftarrow\mathcal{AGGREGATE}_{k}\{\mathcal{GAT}(h^{k-1}_{u^{\prime}},W^{k})\,\forall u^{\prime}\in N_{k}(u)\}$;
@@
$z_{u}\leftarrow\mathcal{COMBINE}(h_{u}^{i}),\forall i\in range(1,K+1)$
@@ 4.2 Distance Recomputator and Topology Reconstructor Implementation
[…]
input features: $x_{v}$, $\forall v\in\mathcal{B}$;
depth $K$: weight matrices $W^{k}$
non-linearity: $\sigma$
differentiable aggregator functions $\text{AGGREGATE}_{k}$, $\forall k\in\{1,...,K\}$
neighborhood sampling functions, $N_{k}:v\longrightarrow 2^{v}$, $\forall k\in\{1,...,K\}$
recompute distance upper-bound: $\alpha$
recompute distance lower-bound: $\beta$
Sample factor: $\gamma$
Init Global NGH = $[0]_{V\times(k\times V)}$
Init Global W = $[\delta_{ij}]_{k\times F}$ //learnable matrix
Output : Vector Representation $z_{v}$ for all $v\in\mathcal{B}$
NGH = K-Hop-Sample($\mathcal{G}$)
@@
[…]
             DRM[i+1][:, :] $\leftarrow$ DRM[i+1][:, :] + DRM[i][DRM[i] ¿ $\alpha$]DRM[i][:, :] $\leftarrow$ DRM[i][:, :] - DRM[i][DRM[i] ¿ $\alpha$]
@@
      for i = 1, …, K do
             DRM[i][:, :] $\leftarrow$ DRM[i][:, :] + DRM[i-1][DRM[i-1] ¡ $\beta$]DRM[i-1][:, :] $\leftarrow$ DRM[i-1][:, :] - DRM[i-1][DRM[i-1] ¡ $\beta$]
@@
Algorithm 3 Distance Recomputator and Asynchronous Aggregator Implementation
@@
Application and Efficiency: The proposed algorithm is not only theoretically robust but also demonstrates practical efficiency in handling GNN tasks. By balancing the computational demands (time complexity) and model performance, our approach offers a feasible solution for real-world applications requiring in-depth graph analysis and dynamic topology reconstruction.
@@ 4.3 Experimental Analysis
Performance Overview: The results, as presented in Table LABEL:tab:widgets, showcase the efficacy of our models in comparison with established GNN methods like GAT, GraphSAGE, GCN, SGC, SSGC, and APPNP. Notably, our models - particularly when integrated with existing GNN structures (GCN, SGC, SSGC, and APPNP) - demonstrate superior performance across all datasets.

### v2

@@ 4.1 In-Depth Learning in Graph Neural Networks
for $n\in\mathcal{V}$ do
       for $k=K\dots 1$ do
             $\mathcal{SN}=\mathcal{RS}(N,\text{Layer}^{(k-1)})$
             for $v_{i}\in\text{Layer}(n,k-1)$ do
                   $\mathcal{NGH}^{(k)}_{n}\leftarrow\mathcal{NGH}^{(k)}_{n}\cup\text{GraphSAGE}(v_{i},\mathcal{SN}[i])$
@@
$\mathcal{B}^{(K)}\leftarrow\mathcal{B}$ // $\mathcal{B}$: Batch of nodes to process
for $k=K\dots 1$ do
       $\mathcal{B}^{(k-1)}\leftarrow\mathcal{B}^{(k)}$
       for $u_{i}\in\mathcal{B}$ do
             $\mathcal{SN}_{s}\leftarrow\mathcal{RS}(u_{i},N)$
             $\mathcal{NGH}^{(k)}_{n}\leftarrow\mathcal{NGH}^{(k)}_{n}\cup\text{GraphSAGE}(u_{i},\mathcal{SN}_{s}[i])$
             $z_{u_{i}}\leftarrow\mathcal{GKHDA}(u_{i},k,\mathcal{NGH}^{(k)}_{n})$
@@
Algorithm 1 Static Sampling and Dynamic Resampling
DRTR Diffusion Layer design is an integration of K-hop diffusion, adaptive message passing, and dynamic topology reconstruction to capture both local and long-range dependencies. Instead of standard 1-hop aggregation, DRTR employs an iterative K-hop aggregation process, where node representations are refined using a combination of local neighborhood aggregation and global diffusion attention. To further improve feature propagation, a Graph Attention Mechanism (GAT) dynamically computes attention weights, prioritizing important neighbors and filtering out noisy connections. Additionally, DRTR introduces a Topology Reconstructor Module, which adjusts edge weights and graph connectivity by recomputing distances and adapting to evolving structures. This mechanism mitigates over-smoothing and ensures the model remains responsive to structural variations. By balancing computational efficiency and expressive power, DRTR’s hierarchical multi-hop framework enables effective large-scale graph learning while maintaining a trade-off between depth (global context) and locality (fine-grained details).
Input :  Graph $\mathcal{G(V,E)}$, Depth $K$, Non-linearity $\sigma$,
Input features $\mathbf{x_{v}}$, Weight matrices $\mathbf{W^{k}}$,
Aggregator functions $\mathcal{AGGREGATE}_{k}$, $\forall k\in\{1,...,K\}$,
Neighborhood sampling $N_{k}:v\to 2^{v}$, $\forall k\in\{1,...,K\}$,
$\mathcal{NGH}\leftarrow\text{StaticSampling}(\mathcal{G},K,\mathbf{Z},\mathbf{A})$
Output : Vector Representation $\mathbf{z_{v}}$ for all $v\in\mathcal{B}$
// Static Aggregation
$h_{u}^{0}\leftarrow x_{v}$, $\forall v\in\mathcal{B}$
for $k=1\dots K$ do
       for $u\in\mathcal{B}^{(k)}$ do
             $h^{k}_{u}\leftarrow\sigma(\mathbf{W}^{k}\cdot(h_{u}^{k-1},h_{N(u)}^{k}))$
             $h^{k}_{u}\leftarrow\frac{h^{k}_{u}}{\lVert h^{k}_{u}\rVert_{2}}$
@@
// K-hop Diffusion Attention
for $k=K\dots 1$ do
       for $u\in\mathcal{NGH}^{(k)}_{u}$ do
             $h^{k}_{N(u)}\leftarrow\mathcal{AGGREGATE}_{k}\{\mathcal{GAT}(h^{k-1}_{u^{\prime}},\mathbf{W}^{k})\,|\,u^{\prime}\in N_{k}(u)\}$
@@
$\mathbf{z_{u}}\leftarrow\mathcal{COMBINE}(h_{u}^{i}),\forall i\in\{1,\dots,K\}$
@@ 4.2 Distance Recomputator and Topology Reconstructor Implementation
Input :  Graph $\mathcal{G(V,E)}$, Depth $K$, Non-linearity $\sigma$,
Input features $\mathbf{x_{v}}$, Weight matrices $\mathbf{W^{k}}$,
Aggregator functions $\mathcal{AGGREGATE}_{k}$, $\forall k\in\{1,...,K\}$,
Neighborhood sampling $N_{k}:v\to 2^{v}$, $\forall k\in\{1,...,K\}$,
Distance bounds $\alpha,\beta$, Sampling factor $\gamma$,
$\mathcal{NGH}\leftarrow[0]_{V\times(K\times V)}$, $\mathbf{W}\leftarrow[\delta_{ij}]_{K\times F}$
Output : Vector Representation $\mathbf{z_{v}}$ for all $v\in\mathcal{B}$
// K-hop Sampling
$\mathcal{NGH}\leftarrow\text{K-Hop-Sample}(\mathcal{G})$
@@
       $\mathcal{NGH}\leftarrow\mathcal{NGH}\cup\text{Resample}(\mathcal{G},\text{ResampleNum})$
       $\mathcal{NGH}_{\text{compute}}\leftarrow\text{Sample}(\mathcal{NGH},\gamma)$
       $\mathbf{Z_{v}}\leftarrow\text{GKHDDRA}(\mathcal{G},\mathcal{NGH}_{\text{compute}})$
       $\mathbf{DRM}\leftarrow\text{mask}(\mathcal{GAT}(\mathcal{G},\mathbf{Z_{v}}),\mathcal{NGH})$
      // Distance Recomputing
       for $i=K-1\dots 0$ do
             $\mathbf{DRM}[i+1]\leftarrow\mathbf{DRM}[i+1]+\mathbf{DRM}[i][\mathbf{DRM}[i]>\alpha]$
             $\mathbf{DRM}[i]\leftarrow\mathbf{DRM}[i]-\mathbf{DRM}[i][\mathbf{DRM}[i]>\alpha]$
@@
      // Topology Reconstruction
       for $i=1\dots K$ do
             $\mathbf{DRM}[i]\leftarrow\mathbf{DRM}[i]+\mathbf{DRM}[i-1][\mathbf{DRM}[i-1]<\beta]$
             $\mathbf{DRM}[i-1]\leftarrow\mathbf{DRM}[i-1]-\mathbf{DRM}[i-1][\mathbf{DRM}[i-1]<\beta]$
@@
Algorithm 3 Distance Recomputator and Topology Reconstructor
@@ 4.3 Experimental Analysis
Performance Overview: The results, as presented in Table 2, showcase the efficacy of our models in comparison with established GNN methods like GAT, GraphSAGE, GCN, SGC, SSGC, and APPNP. Notably, our models - particularly when integrated with existing GNN structures (GCN, SGC, SSGC, and APPNP) - demonstrate superior performance across all datasets.